Company produces three products: Product A, Product B, and Product C.

           During the year the joint costs of processing the three products were $400,000. Production and sales value information were as follows:

                                             Sales Value

Product               Units         at Split-Off        Separable Costs       Selling Price

A                       400,000      $20 per unit           $ 20 per unit           $40 per unit

B                       400,000       $18 per unit           $ 15 per unit           $28 per unit

C                       800,000      $12 per unit           $14 per unit            $17 per unit

a.   Allocate the joint costs using the physical output method.

b.   Allocate the joint costs using the net realizable value method.

c. Allocate the joint costs using sales value at split-off point method.

CLEAR TABLES_PHOTOS NOT ALLOWED
Answer #1
a.   Allocate the joint costs using the physical output method.
Product Output units Weight Joint Cost
a b c=b/1600000 d=c*400000
A 400000 0.25 $            1,00,000
B 400000 0.25 $            1,00,000
C 800000 0.5 $            2,00,000
Total 1600000 1 $            4,00,000
b.   Allocate the joint costs using the net realizable value method.
Product Output units Selling Price Separable Costs Net Realisable Value Weight Joint Cost
a b c d e=(c-d)*b f=e/15600000 g=f*4000000
A 400000 $                                     40 20 $                        80,00,000 0.513 $       2,05,128
B 400000 28 15 5200000 0.333 $       1,33,333
C 800000 17 14 2400000 0.154 $           61,538
Total 1600000 15600000 $       4,00,000
c. Allocate the joint costs using sales value at split-off point method.
Product Output units Sale Price At Split off Sale Value Weight Joint Cost
a b c d=c*b e=d/24800000 f=e*400000
A 400000 20 $          80,00,000 0.323 $                    1,29,032
B 400000 18 $          72,00,000 0.290 $                    1,16,129
C 800000 12 $          96,00,000 0.387 $                    1,54,839
Total 1600000 $      2,48,00,000 1 $                    4,00,000